A satellite moves in an elliptical orbit with the Earth at one focus. At the perigee its
speed is v and its distance from the centre of the Earth is R . If the eccentricity of its
orbit is 0.5, calculate its speed at the apogee.
1
Expert's answer
2019-04-15T10:15:36-0400

Let "v_a" denote the speed of the satellite at the apogee to be found, and let its distance from the center of the Earth at the apogee be "R_a". The conservation of energy of the satellite gives the equation

"\frac{m v_a^2}{2} - \frac{G m M}{R_a} = \frac{m v^2}{2} - \frac{G m M}{R} \, ,"

where "m" and "M" are the masses of the satellite and of the Earth, respectively, and "G" is Newton's gravitational constant. Solving this equation with respect to "v_a", we have

"v_a = \sqrt{ v^2 - \frac{G M}{R} \left( 1 - \frac{R}{R_a} \right) } \, ."

The ratio of the distances in the perigee and apogee is related to the eccentricity "\epsilon" as

"\frac{R}{R_a} = \frac{1 - \epsilon}{1 + \epsilon} \, ."

Since "\epsilon = 1/2", we obtain "R \/ R_a = 1\/3" and

"v_a = \sqrt{ v^2 - \frac{2 G M}{3 R} } \, ."



Answer: "v_a = \sqrt{ v^2 - \frac{2 G M}{3 R} } \, ."